# Incorporating Control Inputs in the Estimation of Continuous Mobile Robot Trajectories and Continuum Robot Shapes

### v1

@@ Abstract
Continuous-time batch state estimation using Gaussian processes is an efficient approach to estimate the trajectories of robots over time. In the past, relatively simple physics-motivated priors have been considered for such approaches, using assumptions such as constant velocity or acceleration. This paper presents an approach to incorporating exogenous control inputs, such as velocity or acceleration commands, into the continuous Gaussian process state-estimation framework. It is shown that this approach generalizes across different domains in robotics, making it applicable to both the estimation of continuous-time trajectories for mobile robots and continuum-robot shapes. Results show that incorporating control inputs leads to more informed priors, potentially requiring less measurements and estimation nodes to obtain accurate estimates. This makes the approach particularly useful in situations in which limited sensing is available.
I Introduction
@@
Throughout this manuscript, we are specifically focusing on continuous-time batch state estimation using Gaussian processes (GPs) to represent the robot’s state and uncertainty continuously over time [2]. This approach involves formulating a suitable prior distribution of continuous robot trajectories. Current formulations are usually based on simple physics models, making specific assumptions about the motion of the robot. In [3] a white-noise-on-acceleration (WNOA) prior is employed, assuming robot motion with constant velocity. This approach has been extended to a white-noise-on-jerk (WNOJ) prior in [4], which assumes constant acceleration. These formulations have been shown to also be applicable to the continuous estimation of the shape of continuum robots [5, 6, 7], substituting time with arclength and velocity with strain. Fig. 1 highlights this analogy. In many cases, the GP prior formulations represent the mobile robot motions or continuum robot shapes well. However, besides the assumptions of constant velocity or acceleration, no additional knowledge about the robot state is incorporated into the prior, such as exogenous control inputs to the system. For mobile robots, these inputs could be known velocity or acceleration commands, while for continuum robots they could consist of known actuation forces and moments acting on their structure. Incorporating such information directly into the GP formulation can lead to more informed prior distributions, representing the mobile robot trajectory or continuum robot shape more accurately. Ultimately, this can mean that less measurements are required to accurately estimate the continuous robot state, when compared to the traditional prior formulations.
@@ I-B Contributions
Throughout this manuscript, we will derive formulations that consider continuous control inputs in , allowing us to approximate the resulting GP prior in closed form. The proposed approach is applied to both the continuous estimation of mobile robot trajectories and continuum robot shapes, highlighting advantages over the existing GP motion priors. We show that considering control inputs is particularly useful when the available measurement data is sparse.
II Continuous State Estimation Using Gaussian Process Regression with Inputs
In the following, we outline derivations for our new GP prior formulation considering control inputs. We refer the reader to [3, 5] for details on how the resulting prior, together with suitable measurements, is to be used in a maximum a posteriori (MAP) batch estimation scheme.
@@ II Continuous State Estimation Using Gaussian Process Regression with Inputs
II-A Gaussian Process Prior
Throughout the following, we are using the notation established in [22]. We describe the continuous-time robot state with the following time-varying stochastic different equations:
Here, is the robot’s pose, is its body-centric velocity, and is a white-noise GP where is the positive-definite power-spectral density matrix. In contrast to the traditional WNOA equations from [3], we additionally consider velocity inputs and acceleration inputs . The overall robot velocity then consists of both the velocity inputs and a velocity bias term driven by white noise and the acceleration inputs. Both inputs are independent and can in theory be applied simultaneously. However, in practice, it is expected we will consider either velocity or acceleration inputs to the system. In the case in which both inputs are zero, the equations simplify to the traditional WNOA case [3].
@@ II-A Gaussian Process Prior
where is the left Jacobian of and inputs , , as well as noise are now defined on the local time interval. Considering the local Markovian state , we can write the following linear time-varying stochastic differential equations:
with
Here, we used the approximation , while also assuming that and are small. The above expression can now be integrated stochastically in closed form to obtain a GP prior (i.e., kernel function) with mean and covariance :
Here, the transition function between two times is the solution to the differential equation
Additionally, is the covariance accumulated between two times,
II-B Prior Error Term
The motion-prior error term for GP regression is
@@ II-B Prior Error Term
We can further express this error term using our global variables as
@@
with . Each of the prior cost terms is a binary factor between two discrete states if thought of as a factor-graph representation. It expresses how close the two corresponding states are to the assumption of the motion prior, considering the incorporated control inputs. We can use these cost terms in an MAP estimation approach, where they are weighed off against any noisy state observations and measurements to obtain a GP posterior of the state (see [3] and [5] for details). It is noted that during an MAP optimization, the integral terms in (10) and (13) only have to be computed during the first iteration and can be held constant afterwards, as they do not depend on the state variables themselves. This significantly reduces the computational burden of the newly introduced derivations.
II-C Gaussian Process Interpolation
@@
Now, we can solve for and considering the mapping between global and local pose variables in (2) and (3). The covariance of the posterior GP can be interpolated similarly. Derivations are omitted for the sake of space.
@@
II-D Example of Piecewise-Linear Inputs
@@
where . Given these expressions for our inputs, we can now find the piecewise transition function between two times and for each individual segment. Given linear inputs in the form of (19) and (20), the system matrix in (9) can be written as a linear function itself such that
with . While there exists no closed form solution to this differential equation, we can use the Magnus expansion [23] to approximate the piecewise transition function for each segment . Keeping the first three terms of this expansion, we have
where is the usual Lie bracket. The total resulting transfer function between the two discrete times and can now be computed by concatenating the individual piecewise ones such that
@@
Here, we are computing the whole integral by splitting it up into several smaller intervals, defined by the discretization of the piecewise-linear input as shown in Fig. 2. By writing out the transition function in these integrals as a product of the individual transition functions of each segment, we see that only one of those terms depends on the integration variable . Given the analytical approximation from (22) and the expression for the piecewise-linear inputs in (19) and (20), we can now solve the integral in closed form. The same strategy can be used to compute the integral for in (10), which we need for the prior covariance in (8). Given the above derivations, further simplifications can be made if the inputs between two times are constant, in which case , or zero, in which case the derivations will simplify to the constant-velocity prior discussed in [3].
@@
As mentioned previously, the transition functions and integrals discussed above generally only have to be computed once. Afterwards, they can be kept constant during the MAP optimization, as they do not depend on the state variables themselves. Nevertheless, when querying the posterior GP after convergence at arbitrary times, some transition functions and integrals have to be re-evaluated. However, by storing most of the intermediate results of the initial computations, some terms can be reused for this interpolation, which leads to an increased efficiency. This is particularly useful if the interpolation times coincide with the piecewise-linear discretization shown in Fig. 2.
@@
II-E Examples
The prior of the first example uses piecewise-constant angular velocity inputs featuring discontinuities, leading to a prior shape composed of three concatenated constant-curvature arcs. The second example considers a sinusoidal angular acceleration input, approximated via piecewise-linear inputs, leading to the depicted sinusoidal velocity profile. For both examples, we show the employed prior as well as the resulting posterior, when an additional position measurement at the end of the trajectory is considered. It can be seen that after fusing with the measurement, the posterior maintains the local shape of the prior in both cases. Further, the robot’s velocity and pose can be accurately recovered along the continuous trajectory using GP interpolation. Thus, even though both examples consider the same position measurement, vastly different posterior trajectories are achieved. We further note that the posterior is able to maintain the discontinuities in the velocity profile of the first example. This would not be possible when using the conventional WNOA prior, while considering the velocity inputs as measurements, as it forces smooth and continuous velocity profiles.
III Mobile Robot Trajectory Estimation
III-A Dataset
We use the mobile robot dataset reported in [2]. In this dataset, a mobile robot is driving within an indoor, planar environment featuring 17 tubular landmarks. The robot is equipped with a laser rangefinder, measuring the distance to each landmark at 10 Hz. Using wheel odometry readings, measurements of the robot’s body velocity are available at 10 Hz as well. Ground-truth for the robot trajectory is provided by a Vicon motion capture system.
III-B Localization Evaluation
For evaluation, we consider the localization problem, where landmark locations are known. We are using the wheel odometry readings as piecewise-linear velocity inputs at a rate of 10 Hz to create a GP motion prior. Landmark distance readings are included as discrete measurements, which are fused with the prior to create a posterior estimate. We are particularly interested in how the proposed approach performs compared to the conventional WNOA GP prior baseline, in which odometry readings are considered as velocity measurements of the state, instead of as inputs. For both methods, we optimize the hyperparameters of the prior, namely , on a subset of the available mobile robot driving data. For the WNOA method, we additionally optimize the covariance of the odometry velocity measurements.

### v2

# Sven Lilge and Timothy D. Barfoot Incorporating Control Inputs in Continuous-Time Gaussian Process State Estimation for Robotics

@@ Abstract
Continuous-time batch state estimation using Gaussian processes is an efficient approach to estimate the trajectories of robots over time. In the past, relatively simple physics-motivated priors have been considered for such approaches, using assumptions such as constant velocity or acceleration. This paper presents an approach to incorporating exogenous control inputs, such as velocity or acceleration commands, into the continuous Gaussian process state-estimation framework. It is shown that this approach generalizes across different domains in robotics, making it applicable to both the estimation of continuous-time trajectories for mobile robots and the estimation of quasi-static continuum robot shapes. Results show that incorporating control inputs leads to more informed priors, potentially requiring less measurements and estimation nodes to obtain accurate estimates. This makes the approach particularly useful in situations in which limited sensing is available.
1 Introduction
@@
Throughout this manuscript, we are specifically focusing on continuous-time batch state estimation using Gaussian processes (GPs) to represent the robot’s state and uncertainty continuously over time [2]. This approach involves formulating a suitable prior distribution of continuous robot trajectories. Current formulations are usually based on simple physics models, making specific assumptions about the motion of the robot. In [3] a white-noise-on-acceleration (WNOA) prior is employed, assuming robot motion with constant velocity. This approach has been extended to a white-noise-on-jerk (WNOJ) prior in [4], which assumes constant acceleration. These formulations have been shown to also be applicable to the continuous estimation of the shape of continuum robots [5, 6, 7], substituting time with arclength and velocity with strain. Fig. 1 highlights this analogy. For mobile robots, we seek to estimate their pose and velocity along a continuous trajectory, where the continuous variable is time. For continuum robots, an analogous estimation problem can be formulated, in which we seek to estimate the robot’s pose and strain along their continuous shape. The continuous variable is now arclength and instead of estimating a trajectory over time, we are estimating a quasi-static shape in space. In both cases, the robot’s pose and their derivative are estimated along a continuous smooth curve (trajectory or shape). In many cases, the GP prior formulations represent the mobile robot motions or continuum robot shapes well. However, besides the assumptions of constant velocity (or strain) or acceleration (equivalent to the strain derivative for continuum robots), no additional knowledge about the robot state is incorporated into the prior, such as exogenous control inputs to the system. For mobile robots, these inputs could be known velocity or acceleration commands, while for continuum robots they could consist of known actuation forces and moments acting on their structure. Incorporating such information directly into the GP formulation can lead to more informed prior distributions, representing the mobile robot trajectory or continuum robot shape more accurately. Ultimately, this can mean that less measurements are required to accurately estimate the continuous robot state, when compared to the traditional prior formulations.
@@ 1.2 Contributions
Throughout this manuscript, we will derive formulations that consider continuous control inputs in , allowing us to approximate the resulting GP prior in closed form. The proposed approach is applied to both the estimation of continuous mobile robot trajectories and continuum robot shapes, highlighting advantages over the existing GP motion priors. We show that considering control inputs is particularly useful when the available measurement data is sparse.
2 Continuous State Estimation Using Gaussian Process Regression with Inputs
In the following, we outline derivations for our new GP prior formulation considering control inputs. We further briefly discuss how the resulting prior can be used in a maximum a posteriori (MAP) batch estimation scheme, but refer the reader to [3, 5] for additional details.
@@ 2 Continuous State Estimation Using Gaussian Process Regression with Inputs
2.1 Gaussian Process Prior
Throughout the following, we are using the notation established in [22]. We describe the continuous-time mobile robot state with the following time-varying stochastic differential equations:
These equations describe how the robot’s pose and body-centric velocity evolve over time. In contrast to the traditional WNOA equations from [3], we consider two additional inputs terms, one for velocity and one for acceleration inputs . The overall robot body-centric velocity now consists of both the velocity inputs and a velocity bias term . This bias term describes motion that is not a direct result of the velocity inputs. Its derivative is driven by a white-noise GP with positive-definite power-spectral density matrix . In addition, we consider potential acceleration inputs as part of this derivative.
Given the stated equations, both inputs and are independent from each other (i.e., is not the derivative of ). In theory, both inputs can be applied simultaneously, resulting in robot motion described by a superposition of velocity and acceleration inputs. In practice, we expect to use either velocity or acceleration inputs, with either or set to zero. When both inputs are zero, the equations simplify to the traditional WNOA case discussed in [3].
@@ 2.1 Gaussian Process Prior
where is the left Jacobian of [22] and inputs , , as well as noise are now defined on the local time interval. Considering the local Markovian state , we can write the following time-varying stochastic differential equations:
We can further simplify this expression as
while using the approximations and , and assuming that and are small. The above expression can now be integrated stochastically in closed form to obtain a GP prior (i.e., kernel function) with mean and covariance :
Here, is the transition function between two times. It is computed by solving the initial value problem of the differential equation
Additionally, is the covariance accumulated between two times, which can be computed as
2.2 Prior Error Term
Next, we derive an error term, evaluating how well a given state aligns with the derived GP prior. The general motion-prior error term for GP regression is
@@ 2.2 Prior Error Term
Using (2) and (3), we can further express this error term using our global variables as
@@
with . Each of the prior cost terms is a binary factor between two discrete states if thought of as a factor-graph representation. It expresses how close the two corresponding states are to the assumption of the motion prior, considering the incorporated control inputs.
2.3 Maximum A Posteriori Batch State Estimation
We can use these cost terms in a MAP batch state estimation approach, where they are weighed off against any noisy state observations and measurements. This allows us to obtain a GP posterior of the state at discrete times . We can again use (2) and (3) to obtain a posterior estimate of our global state variables and at these times. We refer to [3] and [5] for details on the implementation of the MAP batch state estimation approach.
It is noted that during the MAP optimization, the transition function in (9), the accumulated covariance in (10), and the integral term in (13) only have to be computed during the first iteration and can be held constant afterwards, as they do not depend on the state variables themselves. This significantly reduces the computational burden of the newly introduced derivations.
2.4 Gaussian Process Interpolation
@@
Now, we can solve for and considering the mapping between global and local pose variables in (2) and (3). The covariance of the posterior GP can be interpolated similarly. Derivations are omitted for the sake of space, but details can be found in [22, Sec. 11.3] and [23]. This interpolation scheme can further be used during the MAP optimization to incorporate measurement terms at query times different from our discrete times [15].
Although querying the trajectory remains , it may be computationally more expensive than in the WNOA case [3], depending on the inputs and . This increase arises from two factors: (1) the transition function and accumulated covariance may require more computation, and (2) the integral terms in (18) add additional overhead.
@@
2.5 Example of Piecewise-Linear Inputs
@@
where . Given these expressions for our inputs, we can now find the piecewise transition function between two times and for each individual segment. Given linear inputs in the form of (19) and (20), the system matrix in (9) can be written as a linear function itself on the local interval such that
with . Solving this ODE allows us to describe the transition function on the local interval . While there exists no closed-form solution to this differential equation, we can use the Magnus expansion [24] to approximate the piecewise transition function for each segment . The Magnus expansion expresses the result of a linear ODE, such as the one in (21), as an exponential series:
The first three terms of are given as
where is the usual Lie bracket. Because is linear, given our piecewise-linear inputs, we can compute these expressions in closed form. Using these first three terms of the Magnus expansion, we now have
Intuitively, including more terms in the Magnus expansion would improve the approximation of our transition function, but we find the first three terms sufficiently accurate. The total resulting transition function between the two discrete times and can now be computed by concatenating the individual piecewise ones such that
@@
Here, we are computing the whole integral by splitting it up into several smaller intervals, defined by the discretization of the piecewise-linear input as shown in Fig. 2. By writing out the transition function in these integrals as a product of the individual transition functions of each segment, we see that only one of those terms depends on the integration variable , allowing us to pull the remaining terms out of the integral. Given the analytical approximation from (26) and the expression for the piecewise-linear inputs in (19) and (20), we can now solve the integral in closed form. We omit the details of this derivation for brevity. We further note that we truncate the computed integral after the fifth-order terms, which we find sufficiently accurate for implementation.
The same strategy can be used to compute the integral for the accumulated covariance in (10), which we need for the prior covariance in (8):
Using the same strategy as above, we see that several constant terms can again be pulled out of the integral and we can solve the remaining expression using the analytical approximation from (26). We again omit the detailed steps for solving this integral and truncate the result after the fifth order during implementation, which we find sufficiently accurate.
Given the above derivations, further simplifications can be made if the inputs between two times are constant, in which case , or zero, in which case the derivations will simplify to the constant-velocity prior discussed in [3].
@@
As mentioned previously, the transition function, accumulated covariances and integrals discussed above generally only have to be computed once. Afterwards, they can be kept constant during the MAP optimization, as they do not depend on the state variables themselves. Nevertheless, when querying the posterior GP after convergence at arbitrary times, some transition functions, accumulated covariances, and integrals must be re-evaluated, which is now more computationally expensive than in the WNOA case [3]. However, by storing most of the intermediate results of the initial computations, some terms can be reused for this interpolation, which leads to an increased efficiency. This is particularly useful if the interpolation times coincide with the piecewise-linear discretization shown in Fig. 2.
2.6 Examples
The prior of the first example uses piecewise-constant angular velocity inputs featuring discontinuities, leading to a prior shape composed of three concatenated constant-curvature arcs. The second example considers a sinusoidal angular acceleration input, approximated via piecewise-linear inputs, leading to the depicted sinusoidal velocity profile. For both examples, we show the employed prior as well as the resulting posterior, when an additional position measurement at the end of the trajectory is considered. It can be seen that after fusing with the measurement, the posterior maintains the local shape of the prior in both cases. Further, the robot’s velocity and pose can be accurately recovered along the continuous trajectory using GP interpolation. Thus, even though both examples consider the same position measurement, vastly different posterior trajectories are achieved. We further note that the posterior is able to maintain the discontinuities in the velocity profile of the first example. This would not be possible when using the conventional WNOA prior, which could consider the velocity inputs as measurements, as it forces smooth and continuous velocity profiles.
3 Mobile Robot Trajectory Estimation
3.1 Dataset
We use a mobile robot dataset similar to the one reported in [2]. In our dataset, a mobile robot is driving within an indoor, planar environment featuring 17 tubular landmarks. The overall robot trajectory includes 21 minutes of driving. The robot is equipped with a laser rangefinder, measuring the distance to each landmark at 10 Hz. Using wheel odometry readings, measurements of the robot’s body velocity are available at 10 Hz as well. Ground-truth for the robot trajectory is provided by a Vicon motion capture system.
3.2 Localization Evaluation
For evaluation, we consider the localization problem, where landmark locations are known. We are using the wheel odometry readings as piecewise-linear velocity inputs at a rate of 10 Hz to create a GP motion prior. Landmark distance readings are included as discrete measurements, which are fused with the prior to create a posterior estimate. We employ a MAP estimation scheme that computes the posterior distribution of the entire trajectory in a single batch. We are particularly interested in how the proposed approach performs compared to the conventional WNOA GP prior baseline, in which odometry readings are considered as velocity measurements of the state, instead of as inputs. For both methods, we optimize the hyperparameters of the prior, namely , on a subset of the available mobile robot driving data. For the WNOA method, we additionally optimize the covariance of the odometry velocity measurements. The covariance for the landmark distance measurements is given.
@@ 3.2 Localization Evaluation
3.3 Results
The results are summarized in Tab. 1, including the estimation errors and computation times for increasing times between landmark distance measurements . When including a discrete state at every odometry update, both methods achieve similar results. The resulting accuracies differ only by a few mm and less than half of a degree, while the WNOA GP prior baseline achieves shorter computation times. In each case shown, the 21-minute trajectory can be estimated in just a few seconds. When including discrete states only when landmark distance measurements occur, the proposed prior formulation using inputs significantly outperforms the baseline. While the traditional WNOA GP prior is able to achieve competitive accuracies for short time distances , it fails to accurately represent the local shape of the trajectory when landmark measurements occur less frequently ( s). The proposed GP prior with inputs is able to maintain the estimation accuracy for longer times. Here the error only starts to significantly increase for s. Both methods show significantly reduced computation times when excluding states without landmark distance measurements in the batch estimation. However, the computation time increases when using a prior with inputs as increases. In these cases, more states have to be recovered using the GP interpolation schemes, which requires additional effort for the proposed method to solve the Magnus expansion and integral terms.
We conclude that using the proposed GP prior with inputs presents an interesting use case, as it is able to maintain the estimation accuracy, while significantly reducing the computational effort. This is thanks to its ability to maintain the local shape of the odometry over longer time intervals without needing to include discrete states in the batch estimation. We can exploit this advantage, as long as does not grow too large, in which case the accuracy and computation times are not competitive with the WNOA GP prior baseline, even without excluding discrete states from the batch estimation. In our case, it is beneficial to use the newly proposed method for s, excluding states without landmark distance measurements. For s the original WNOA method should be applied, including all discrete states and treating odometry velocities as measurements.
To further illustrate the differences of the two approaches, Fig. 4 shows state estimation results for the first 25% of the mobile robot dataset, while considering landmark distance measurements every s and including discrete estimation states only at these measurement times. The proposed method is able to accurately represent the local shape of the trajectory between the discrete states. In contrast, the baseline, which relies on the constant-velocity prior, cannot accurately capture the local shape of the trajectory and instead produces a highly smoothed result. This leads to considerable inaccuracies in the resulting state estimation. This is particularly evident in the zoomed-in area of the trajectory, which includes a sharp turn. Here, the robot stopped and pivoted on the spot between two discrete estimation times, resulting in a non-smooth trajectory. This motion is captured by the prior that considers velocities as inputs, but cannot be represented by the WNOA prior, whose interpolation overly smooths the trajectory, deviating from the ground-truth.
4 Continuum Robot Shape Estimation
As discussed in detail in [5], the GP batch state estimation approach from [3] for mobile robot trajectories can also be applied to continuously estimate quasi-static continuum robot shapes. Consequently, the proposed GP prior that incorporates inputs can be extended to this domain as well. Analogous to (1a), we can describe the continuum robot shape along its arclength with the following stochastic differential equation:
These equations describe the evolution of the robot’s pose and strain expressed in the body frame along its arclength. Together, they define the quasi-static shape of the continuum robot. Similar to the mobile robot equations, we consider two potential inputs: an input strain and distributed forces and moments , which can act as an input to the system. These distributed loads are mapped to the strain derivative through the stiffness matrix of the continuum robot’s backbone. Lastly, is again a white-noise GP with positive-definite power-spectral density matrix . Using this expression and the previous derivations, we can derive an analogous GP prior with inputs for the continuous estimation of the quasi-static shape of continuum robots.
Throughout this section, we validate this GP prior formulation with additional continuum robot experiments.
4.1 Experimental Setup
For the experiments, we are using the tendon-driven continuum robot (TDCR) prototype and sensor setup presented in [5] (see Fig. 5, left). This prototype features two controllable segments, with tendons routed parallel to the backbone through spacer disks and attached at the end of each segment. The bending of both segments can be individually actuated by pulling and releasing the tendons. For sensing, we consider a single six-degree-of-freedom pose measurement at the robot’s tip using an electromagnetic tracking sensor (Aurora v3, Northern Digital Inc., Canada). Ground-truth is obtained using a coordinate measurement arm with an attached laser probe (FARO Edge with FARO Laser Line Probe HD, FARO Technologies Inc., USA). This measurement arm allows us to capture a detailed laser scan of the robot’s shape in the form of a point cloud, from which the discrete pose of each spacer disk can be reconstructed. Six simple bending configurations of the continuum robot were exclusively recorded to calibrate the sensor and ground-truth frames. Using the same procedure outlined in [5], the remaining calibration error at the tip of the robot results in mm.
4.2 Evaluation
For evaluation, we consider nine different TDCR configurations, with three sets of applied tendon tensions. Each set of tensions is shared across three configurations. In each configuration, an additional unknown force is applied to the robot’s tip to induce further deflection.
For state estimation, we now wish to compare the conventional prior formulation from [5] with the newly proposed one considering inputs. As discussed above, there are two possibilities to incorporate known inputs. One can either consider known input strains or known distributed actuation loads . Either of these possible outputs can, for instance, be provided by another open-loop continuum robot model.
We use the second approach and convert the known applied tendon tensions to moments acting on the robot backbone, considering the simplified point-moment model discussed in [25]. Using this model, we compute the discrete moments applied to the tendon termination points. We then approximate these discrete moments with piecewise-linear inputs to obtain an expression for , which we map to the strain derivative using the known stiffness of the robot’s Nitinol backbone, based on properties reported by the manufacturer. While this simplified model exhibits inaccuracies in out-of-plane bending scenarios [25], we deem it sufficiently accurate to be fused with additional measurements. Alternatively, more complex actuation load distributions could be incorporated, as long as they can be expressed in the robot’s body frame independently from the state variables.
In the mobile robot experiments, velocity inputs were included as measurements in the traditional WNOA method for a fairer comparison. However, we do not include the inputs as measurements for the baseline method in the continuum robot evaluations. This is because our inputs are defined on the strain derivative, which is not part of the system state, making it difficult to directly include such measurements in a straightforward manner.
Both the traditional prior from [5] as well as the newly proposed prior considering inputs are fused with a single pose measurement at the tip of the robot. As an additional degenerate case, we further consider the scenario in which only the robot tip’s position can be measured. For every case, we compute the RMSE and maximum errors for the robot’s position and orientation along its shape, considering the discrete ground-truth poses at each disk.